The resistance of two forces which are of 130N and 90N is 150N. Calculate the angle between the two forces
1
Expert's answer
2019-03-19T12:19:02-0400


Using the cosine rule,


"|{\bf F_1}-{\bf F_2}|^2=F_1^2+F_2^2-2F_1F_2\cos \theta"

we obtain

"\cos \theta=\frac{F_1^2+F_2^2-|{\bf F_1}-{\bf F_2}|^2}{2F_1F_2}"

"\cos \theta=\frac{130^2+90^2-150^2}{2\times130\times90}=0.1068"

"\theta=84^{\circ}"